OpenStudy (anonymous):

A person standing at the edge of a seaside cliff kicks a stone over the edge with a speed of 19 m/s. The cliff is 68 m above the water’s surface, as shown.

OpenStudy (anonymous):

Is there a question in this? It sounds like you're asking about movement in two dimensions, but there is no actual inquiry or problem to solve.

OpenStudy (anonymous):

68=0.5(9.81)(t^2) t=sqrt(2(68)/9.81) x=19(t) That should give you horizontal distance travelled which is what I assumed you wanted

OpenStudy (anonymous):

If you wanted the horizontal distance, degret is correct.